OpenStudy (anonymous):

Write an equation having the given slope and y intercept. M= 3 over 2 b= -3 Medals!

OpenStudy (anonymous):

y=mx+c m=slope=3/2 x=x-intercept c= y-intercept=-3 put the values in the equation y=3/2x+(-3) solve it more :)